OpenStudy (anonymous):

verify the identity and show all work 1- tanxtany= cos(x+y)/cosxcosy

hartnn (hartnn):

write tan in terms of sin and cos

OpenStudy (anonymous):

so 1-(sinx/cosx)sin x/sin y

hartnn (hartnn):

tan y = ?

OpenStudy (anonymous):

siny/cosy

hartnn (hartnn):

so it would be \(\huge 1-\frac{sin x}{cosx}\times \frac{sin y}{cos y}\) right ? now cross-multiply.

OpenStudy (anonymous):

1-sinxcosy sinycosx

hartnn (hartnn):

it would be \(\huge \frac{cos x cos y -sin x sin y}{cos x cos y}\) got this ?

OpenStudy (anonymous):

ok yea

hartnn (hartnn):

now numerator is the standard formula for cos (x+y)

OpenStudy (anonymous):

ohh ok got it !(: